OpenStudy (anonymous):

Does the graph represent the same equation as the table? Explain. A graph of the following ordered pairs: zero comma three, one comma five, two comma seven, and three comma nine. A table of x- and y-values. When x is 4 y is 5, when x is 5 y is 7, and when x is 6 y is 9. im i correct? no its dosnet because 4,5 5,7 6,9 is not in graph only 0,3 1,5 2,7 and 3,9 are only in the graph so no the graph does not represent the same equation as the table.

OpenStudy (anonymous):

the graph...\[\begin{array}{|c|c|c|c|c|} \hline x&0&0&2&3\\ \hline y&3&5&7&9\\ \hline \end{array}\]the table... \[\begin{array}{|c|c|c|c|} \hline x&4&5&6\\ \hline y&5&7&9\\ \hline \end{array}\]

OpenStudy (anonymous):

analyzing the data you provided for both graph and table...

OpenStudy (anonymous):

the data on the table represents a line with a slope of 2, taking (4,5) as a sample point...\[(y-y_1)=m(x-x_1)\]\[(y-5)=2(x-4)\]\[y-5=2x-8\]\[y=2x-8+5=2x-3\]\[y=2x-3\]testing if the data on the graph satisfy the equation obtain from the table... check for (3,9)...\[is~9=2(3)-3=6-3=3~?\]...obviously not. \[\LARGE \sf \color{#FF2E00}{Therefore~ it~ does~ not.}\]